OpenStudy (anonymous):

I need someone who would know how to convert a log to a base of 10? I'll post the question and what work I have done in the comments :)

OpenStudy (anonymous):

The computer system you need to input the function into only works in logarithms of base 10. Using complete sentences, explain how to convert your exponential function P(x) in a logarithmic one and then into a base 10 logarithm. Log10 / log2 --> (1) / (0.301) = 3.321 not sure if this is correct or not but the equation I used was P(x) = 2^2 --> log_2 4 = 2

OpenStudy (imstuck):

What is your exponential function? Can you give me the problem as it is in your book or whatever?

OpenStudy (anonymous):

Yes hold on :)

OpenStudy (anonymous):

Months (x) Population P(x) 0 2 1 4 2 16 3 256 x^2 = P(x)

OpenStudy (anonymous):

This is what I did and numbers I plugged in @IMStuck

OpenStudy (imstuck):

will have to come back to you; my family needs me to cook dinner now! But I'll be back in a bit! Promise!

OpenStudy (anonymous):

Oh okay :( Well thanks and I hope to get help when you get back :)

OpenStudy (jdoe0001):

\(\Large \begin{array}{llll} x\qquad &y \\\hline\\ {\color{brown}{ 0}}&2\to 2^1\to 2^{2^{\color{brown}{ 0}}}\\ {\color{brown}{ 1}}&4\to 2^2\to 2^{2^{\color{brown}{ 1}}}\\ {\color{brown}{ 2}}&16\to 2^4\to 2^{2^{\color{brown}{ 2}}}\\ {\color{brown}{ 3}}&256\to 2^8\to 2^{2^{\color{brown}{ 3}}} \end{array}\)

OpenStudy (anonymous):

Are those logs? That confused me a little sorry :(

OpenStudy (jdoe0001):

ahemm... no, just p(x), meaning that \(\huge \bf p(x)= 2^{2^{\color{brown}{ x}}}\)

OpenStudy (anonymous):

so is that my equation? Sorry that makes no sense to me I'm completely confused :S

OpenStudy (jdoe0001):

heehhe

OpenStudy (jdoe0001):

well, you're trying to get a logarithmic equation from the dataset provided, right?

OpenStudy (anonymous):

I have no clue how to get a log out of that so would you mind helping me find it?

OpenStudy (jdoe0001):

\(\Large \bf recall\implies {\color{red}{ a}}^y={\color{blue}{ b}}\implies log_{\color{red}{ a}}{\color{blue}{ b}}=y\)

OpenStudy (anonymous):

so log_2 4 = 2 is my log ?

OpenStudy (jdoe0001):

well \(\bf recall\implies {\color{red}{ a}}^y={\color{blue}{ b}}\implies log_{\color{red}{ a}}{\color{blue}{ b}}=y \\ \quad \\ {\color{blue}{ p(x)}}={\color{red}{ 2}}^{2^x}\implies log_{{\color{red}{ 2}}}{\color{blue}{ p(x)}}=2^x\)

OpenStudy (anonymous):

how do I change the base to 10?

OpenStudy (jdoe0001):

use the "change of base rule", that is \(\bf \textit{log change of base rule }log_{\color{red}{ a}}{\color{blue}{ b}}\implies \cfrac{log_{\color{olive}{ c}}{\color{blue}{ b}}}{log_{\color{olive}{ c}}{\color{red}{ a}}}\qquad thus \\ \quad \\ log_2[p(x)]=2^x\implies \cfrac{log_{{\color{olive}{ 10}}}[p(x)]}{log_{{\color{olive}{ 10}}}2}=2^x\)

OpenStudy (jdoe0001):

for the log change of base rule, the log base can be any number, so long it's the same above and below

OpenStudy (jdoe0001):

in this case we just happen to use base 10